# What Makes a Language Easy to Deep-Learn?

### v3

@@ 1 Introduction
Compositionality, i.e., whether the meaning of a compound expression can be derived solely from the meaning of its constituent parts, has been studied for decades by both computer scientists and linguists [1, 2, 3, 4, 5]. In particular, languages differ in how they map meanings into morphosyntactic structures [6, 7] and cross-linguistic studies find substantial differences in the degree of structural complexity across languages [8, 9, 10, 11, 12, 13, 14]. These differences can stem from multiple and often confounded aspects of linguistic structure including the degree of compositionality [7], which can be quantified by correlating differences in meaning with differences in form [15]. For example, the English term “white horse” is compositional since its meaning can be directly inferred given knowledge about its constituents “white” and “horse”. In contrast, consider the equivalent German term “Schimmel”, whose meaning cannot be derived from “weiß” (white) and “Pferd’ (horse). Crucially, compositionality directly affects our ability to make systematic generalizations in a given language and thus shapes its immense expressive power – which also explains its high relevance in machine learning [2, 16, 17, 18, 19, 20, 21, 22, 1, 23].
@@
Importantly, cross-linguistic differences in compositionality were suggested to impact human language learning and generalization in the real world [24, 25, 26] as well as in lab experiments [27, 28, 29, 30, 31], with more compositional linguistic structures typically being easier to learn for adult learners. In a large-scale artificial language learning study with adult human participants, the acquisition of a broad yet tightly controlled range of comparable languages with different degrees of compositional structure was tested [27]. Results showed that more compositional languages were learned faster, better, and more consistently by the adult learners, and that learning more structured languages also promoted better generalizations and more robust convergence on labels for new, unfamiliar meanings. This is likely because more systematic and compositional linguistic input allow learners to derive a set of generative rules rather than rote memorizing individual forms, and then enables learners to use these rules to produce an infinite number of utterances after exposure to just a finite set [32, 33, 34, 31, 35]. This learnability and generalization advantage for more structured linguistic input has far-reaching implications for broader theories on language evolution in our species (and potentially other learning systems): A large body of computational models and experimental work with human participants show that more systematic and compositional structures emerge during cross-generational transmission and communication precisely because such structures are learned better, while still allowing for high expressivity [32, 33, 34, 31, 36, 29, 37, 38, 39]. Hence, popular theories of language evolution attribute the emergence of systematic and compositional structure in natural languages to such learnability pressures [31, 40], suggesting a a causal role not only in language learning, but also in shaping the way human languages are structured. To what extent this advantage of linguistic structure carries over to artificial learning systems is currently poorly understood – which is the aim of the current study.
@@
Despite the increasing body of work that reports striking similarities between humans and large language models [41, 42, 43, 44, 45, 46, 47], and despite large language models being incredibly proficient at using language and generalizing to new tasks with little to no new training data [48, 49, 50, 51], research on emergent communication suggests that deep neural networks (the class of models that underlies large language models) show no correlation between the degree of compositional structure in the emergent language and the generalization capabilities of the networks. In other words, unlike humans, artificial neural networks do not seem to benefit from more compositional structure when they are made to develop their own communication protocol, at least without dedicated intervention [52, 53, 54, 55]. Thus, this finding raises the question of whether systematic and compositional linguistic structure is helpful at all for deep neural networks, and to what extent does compositionality affect the memorization and generalization abilities of deep neural networks learning a new language.
@@
Figure 1: Overview of input languages (Top), the experimental procedure (Left), and model architecture (Right). Low-structured input languages show no signs of systematicity or compositionality, whereas high-structured languages are systematic and compositional with respect to both attributes: shape and angle. For each language, we train the model for multiple rounds of exposure, guessing, production. After each round, we conduct a memorization test to evaluate productions for previously seen items and a generalization test evaluating the productions for new items.
@@
To allow for direct comparisons between humans and machines, we carefully follow the experimental procedure and measures of a recent large-scale preregistered language learning study with adult participants [27]. Figure 1 shows the recurrent neural network architecture and summarizes the experimental procedure: Neural networks were trained on the exact same stimuli presented to humans and in the same order, using the same learning tasks, providing the same feedback during learning blocks, and evaluated with the same memorization and generalization tests. Full details of the experimental setup, custom recurrent neural neural network models, and how we employed large language models are provided in the Methods section.
2 Results
To preview our results, we find a consistent advantage of more systematic and compositional linguistic structure for learning and generalization, closely reflecting adult human participants. The generalization behavior of both large language models (pre-trained on other languages) and recurrent neural networks (trained from scratch) was far more systematic and transparent when the input languages were more compositional. Moreover, recurrent neural network agents displayed a higher agreement with other agents as well as with humans when the input was more compositional, leading to converging transparent generalizations for new unseen input.
@@ 2 Results
Figure 2: Final Generalization Score achieved by humans (A), GPT-3 (B), and recurrent neural networks (C) for each of the input languages. The x-axis shows the structure score of the input languages. Error regions show 95% confidence intervals estimated via bootstrapping. More structure in the input language leads to more systematic generalization for all three learning systems.
@@
Table 1: Generalization examples from neural network and human learners, showing labels generated for unseen scenes. The column GPT-3 corresponds to completions generated by the GPT-3 model text-davinci-003 via in-context learning, where the training data is provided in context. The examples cover the differently structured input languages from low to high.
Struct. Shape Angle Human RNN GPT-3
@@
Strikingly, the results reveal that a higher degree of compositional structure in the input language leads to generalizations that are more systematic (see Figure 2B), closely reflecting the pattern of adult human learners (Figure 2A). Table 1 shows examples of the final productions of humans and large language models during generalization.
In addition, we evaluate the production similarity as character-level length-normalized edit distance between the generated labels and labels produced by human participants during generalization. The results show that, given more structured linguistic input, GPT-3 also yields productions that are more similar to the productions of human participants, calculated as the average similarity between GPT-3’s production and all human productions for the same scene in the same language (Figure 3A).
Figure 3: Final Similarity to Humans during Generalization Final production similarity with human participants during generalization achieved by GPT-3 (A) and recurrent neural networks (B) for each of the input languages. The x-axis shows the structure score of the input languages. Error regions show 95% confidence intervals estimated via bootstrapping. More structure in the input language leads to more similarity to human participants for both RNNs and GPT-3.
@@
Figure 4: Memorization Error Analysis for human participants (A), GPT-3 (B), and recurrent neural networks (C). The error rates are 33.30% for humans, 7.39% for GPT-3, and 13.87% for RNNs. The x-axis shows the structure score of the input language. Error regions show 95% confidence intervals estimated via bootstrapping. More structure leads to erroneously memorized examples being more similar to the ground truth of the input language.
@@
Figure 5: Learning trajectory of Recurrent Neural Networks’ Memorization and Generalization Performance. More structured languages lead to better and faster reproduction of the input language (a), to better generalization on unknown scenes (c), better agreement with human participants during memorization (b) and generalization (d), and higher convergence between networks (e). (A) Production similarity between labels generated by neural agents and labels of the input language. (B) Production similarity between labels generated by neural agents and labels generated by human participants. (C) Generalization score of labels generated by neural agents for new scenes that were not part of the training data. (D) Production similarity between labels generated by neural agents and labels generated by human participants for unseen scenes. (E) Convergence score measures the similarity between labels generated for unseen scenes by different neural agents. Stars mark the round at which neural agents first exceed the final performance of human participants. Input languages are grouped into 5 bins. Each line is the average of 200 neural agents with a different random initialization. A star marks the epoch at which the RNN agents exceed human performance. Results are cut off for visualization at epoch 60, full results in Appendix B.
@@
The results are shown in Figure 5. Extended results can be found in Appendix B. In the following, we present the results for the learning trajectory organized along the two types of tests: memorization and generalization, before presenting the final results of RNNs.
@@
Similarity to input languages during memorization: With sufficient training rounds, all languages can be learned by all neural network agents, reaching a production similarity of at least 0.8 (out of 1) by round 60 (Figure 5A). Structured languages are learned significantly better (LME 1; $\beta=0.045$, $\mathrm{SD}=0.001$, $z=62.865$, $p<0.001$), i. e., they show a higher similarity with the input language. However, this advantage tends to diminish over training rounds (LME 1; $\beta=-0.005$, $\mathrm{SD}<0.001$, $z=-54.978$, $p<0.001$).
@@
Similarity to Humans during Generalization We measure the similarity between the productions of neural network agents and humans for new scenes (Figure 5D), i. e., during generalization. Examples are shown in Table 1. More structure in the input language led to a significantly higher similarity between humans and neural agents (LME 5; $\beta=0.132$, $\mathrm{SD}=0.002$, $z=70.280$, $p<0.001$), which became stronger over rounds ($\beta=0.046$, $\mathrm{SD}<0.001$, $z=344.287$, $p<0.001$).
Convergence between Neural Agents during Generalization More structured languages lead to better agreement between networks (LME 4; $\beta=0.043$, $\mathrm{SD}=0.001$, $z=49.027$, $p<0.001$), such that, for more structured languages, different neural agents learning the same input language produced more similar labels for new scenes (Figure 5E). This effect became stronger over rounds ($\beta=0.009$, $\mathrm{SD}<0.001$, $z=121.740$, $p<0.001$).
@@
To compare our custom recurrent neural network agents with large language models and with humans, we visualize the relationship between compositional structure of the input language and final generalization performance in Figure 2C. All three learning systems (Humans, RNNs, and GPT-3) show the same trend: more compositionality in the input language leads to more systematic generalization.
@@
Moreover, we calculate the similarity to human participants who received precisely the same learning input in the precise same order. Comparing the productions during generalization, the results show that a higher degree of structure in the input language leads to more similarity with humans (see Figure 3B). This pattern of compositional structure leading to more human-like generalizations is present in both RNNs’ and GPT-3’s generated labels.
@@
Lastly, we visualize the results of the memorization error analysis for recurrent neural networks alongside humans and GPT-3 in Figure 4C. The pattern is the same for all three different learning systems, be it artificial or biological: more compositional structure leads to errors that are more similar to the true label.
3 Discussion
@@
Our results show that deep neural networks benefit from more structured linguistic input as humans do and that neural networks’ performance becomes increasingly more human-like when trained on more structured languages. This structure bias can be found in networks’ learning trajectories and even more so in the networks’ generalization behavior, mimicking previous findings with humans. Although all languages can eventually be learned, we show that more structured languages are learned better and more similarly to human productions. Deep neural networks and humans produce nearly identical labels when trained on high-structured languages but not when trained on low-structured languages. Moreover, networks that learn more structured languages are significantly better at systematic generalization to new, unseen items, and crucially, their generalizations are significantly more consistent and more human-like. This means that highly systematic grammars allow for better generalization and facilitate greater alignment between different neural agents and between neural agents and humans. We have replicated these results with small recurrent neural networks and with transformer-based large language models, showing that, together with humans, all three learning systems show the same bias in systematic generalization and memorization errors. Thus, our findings strengthen the idea that language models are useful for studying human cognitive mechanisms, complementing the increasing evidence of similarity in language learning between humans and machines [41, 42, 43, 44, 45, 46, 47]
@@
First and foremost, our findings support the idea that languages’ underlying grammatical structure can be learned directly from (grounded) linguistic input alone [71, 72, 73, 40, 34]. To ensure that the advantage of more structured linguistic input does not stem from the fact that the learning system was already proficient in a different language – i. e., as are pre-trained language models and adult humans – we also also considered models trained from random initialization. Therefore, our results predict that children would also benefit from more systematic compositional structure in the same way adults do – a prediction we are currently testing (preregistration: [74]).
@@
Our findings have further implications for machine learning, where systematic generalization beyond the training distribution is of high interest [16, 18, 75, 19, 20]. Specifically, we show that seeding a learning system with well-structured inputs can improve their ability to systematically generalize to new inputs. Even though our study is based on artificial languages, our findings directly pertain to the natural language processing of real-world languages. To confirm this prediction, we re-analyzed data from Wu et al. [14], who used the Wug Test [76] to test language models’ ability to predict different forms of unfamiliar words in a wide range of natural languages. Indeed, we find that the Wug Test accuracy negatively correlates with the degree of irregularity of the language (Spearman’s $\rho=-0.96$, $p<10^{-15}$; Kendall’s $\tau=-0.86,p<10^{-14}$). This strong negative correlation suggests that natural languages with fewer irregularities, i. e., more consistently structured natural languages, are indeed easier to learn for machines.
@@
Crucially, there is a positive correlation between the degree of linguistic structure and population size [12, 77, 78, 79], with low-resource languages (i. e., languages spoken by smaller communities for which there is only very little training data available) typically having less structured languages. Since our study predicts that such languages are harder to learn for deep neural networks, this results in a double whammy for developing natural language processing systems for small communities’ languages – exacerbating challenges of developing language models for low-resource languages [80].
@@
In conclusion, our findings shed light on the relationship between language and language-learning systems, showing that linguistic structure is crucial for language learnability and systematic generalization in neural networks. Our results suggest that more structured languages are easier to learn, regardless of the learning system: a human, a recurrent neural network, or a large language model. Thus, generalization capabilities are heavily influenced by compositional structure, with both biological and artificial learning systems benefiting from more structured input by facilitating more systematic and transparent generalizations. In future work, we will analyze how this learnability bias for more structure affects neural networks engaging in collaborative communication games. Moreover, our findings give a clear prediction that children would benefit from more structure in the linguistic input, which we will test by conducting a learnability study with children.
4 Methods

### v4

@@ Abstract
Compositionality, i.e., whether the meaning of a compound expression can be derived solely from the meaning of its constituent parts, has been studied for decades by both computer scientists and linguists [1, 2, 3, 4, 5]. In particular, languages differ in how they map meanings into morphosyntactic structures [6, 7] and cross-linguistic studies find substantial differences in the degree of structural complexity across languages [8, 9, 10, 11, 12, 13, 14]. These differences can stem from multiple and often confounded aspects of linguistic structure including the degree of compositionality [7], which can be quantified by correlating differences in meaning with differences in form [15]. For example, the English term “white horse” is compositional since its meaning can be directly inferred given knowledge about its constituents “white” and “horse”. In contrast, consider the equivalent German term “Schimmel”, whose meaning cannot be derived from “weiß” (white) and “Pferd” (horse). Crucially, compositionality directly affects our ability to make systematic generalizations in a given language and thus shapes its immense expressive power – which also explains its high relevance in machine learning [2, 16, 17, 18, 19, 20, 21, 22, 23, 1, 24].
@@
Importantly, cross-linguistic differences in compositional structure were suggested to impact human language learning and generalization in the real world [25, 26, 27] as well as in lab experiments [28, 29, 30, 31, 32], with more compositional linguistic structures typically being easier to learn for adult learners. In a large-scale artificial language learning study with adult human participants, the acquisition of a broad yet tightly controlled range of comparable languages with different degrees of compositional structure was tested [28]. Results showed that more compositional languages were learned faster, better, and more consistently by the adult learners, and that learning more structured languages also promoted better generalizations and more robust convergence on labels for new, unfamiliar meanings. This is likely because more systematic and compositional linguistic input allow learners to derive a set of generative rules rather than rote memorizing individual forms, and then enables learners to use these rules to produce an infinite number of utterances after exposure to just a finite set [33, 34, 35, 32, 36]. This learnability and generalization advantage for more structured linguistic input has far-reaching implications for broader theories on language evolution in our species (and potentially other learning systems): A large body of computational models and experimental work with human participants show that more systematic and compositional structures emerge during cross-generational transmission and communication precisely because such structures are learned better, while still allowing for high expressivity [33, 34, 35, 32, 37, 30, 38, 39, 40]. Hence, popular theories of language evolution attribute the emergence of systematic and compositional structure in natural languages to such learnability pressures [32, 41], suggesting a causal role not only in language learning, but also in shaping the way human languages are structured. To what extent this advantage of linguistic structure carries over to artificial learning systems is currently poorly understood – which is the aim of the current study.
@@
Despite an increasing body of work that reports striking similarities between humans and large language models [42, 43, 44, 45, 46, 47, 48], and despite large language models being incredibly proficient at using language and generalizing to new tasks with little to no new training data [49, 50, 51, 52], research on emergent communication suggests that deep neural networks (the class of models that underlies large language models) show no correlation between the degree of compositional structure in the emergent language and the generalization capabilities of the networks. In other words, unlike humans, artificial neural networks do not seem to benefit from more compositional structure when they are made to develop their own communication protocol, at least without dedicated intervention [53, 54, 55, 56] (but see [57]). Thus, this finding raises the question of whether systematic and compositional linguistic structure is helpful at all for deep neural networks, and to what extent compositionality affects the memorization and generalization abilities of deep neural networks learning a new language.
@@
Figure 1: Overview of input languages (Top), the experimental procedure (Bottom Center) along with exemplary input data from one language (Bottom Left), and the model architecture (Bottom Right). Low-structured input languages show no signs of systematicity or compositionality, whereas high-structured languages are systematic and compositional with respect to both attributes: shape and angle. For each language, we train the model for multiple rounds of exposure, guessing, production. After each round, we conduct a memorization test to evaluate productions for previously seen items and a generalization test evaluating the productions for new items. Graphical elements in the upper part of this figure are re-used and adapted with permission from Raviv et al. [68].
@@
To allow for direct comparisons between humans and machines, we carefully follow the experimental procedure and measures of a recent large-scale preregistered language learning study with adult participants [28]. We consider 10 input languages, each of which has emerged independently and spontaneously through a group communication experiment with adult human participants [68]. The languages describe four different novel shapes moving on the screen in a different direction (0-360 degree), and vary in their degree of compositional structure: ranging from fully idiosyncratic languages with entirely different labels for two related meaning (e.g., ’kuim’ and ’goom’ for the same shape moving into a different direction) to highly structured languages, which re-use parts of the descriptive label (e.g., referring to the two scenes as ’fest-ii’ and ’fest-ui’). See Figure 1. Neural networks were then trained on the exact same stimuli presented to humans and in the same order, using the same learning tasks, providing the same feedback during learning blocks, and evaluated with the same memorization and generalization tests. Figure 1 shows the recurrent neural network architecture and summarizes the experimental procedure: Full details of the experimental setup, custom recurrent neural neural network models, and how we employed large language models are provided in the Methods section.
Results
To preview our results, we find a consistent advantage of more systematic and compositional linguistic structure for learning and generalization, closely reflecting adult human participants. The generalization behavior of both large language models (pre-trained on other languages) and recurrent neural networks (trained from scratch) was far more systematic and transparent when the input languages were more compositional. Moreover, recurrent neural network agents displayed a higher agreement with other agents as well as with humans when the input was more compositional, leading to converging transparent generalizations for new unseen input. A glossary of evaluation metrics can be found in Table LABEL:tab:glossary. More detailed descriptions of the metrics are provided in the Methods section.
Table 1: Glossary of Metrics
Metric Description
Production Similarity One minus length-normalized edit distance
Semantic Difference Sum of the difference in shape (1 if different and 0 otherwise) and the absolute difference in angles (divided by 180)
Structure Score Pearson correlation between (a) pairwise semantic distances and (b) pairwise normalized Levenshtein distances, where (a) and (b) are calculated on all pairs of items in the original input language
Generalization Score Pearson correlation between (a) pairwise semantic difference and (b) pairwise length-normalized edit distance, where (a) and (b) are calculated on all pairs between productions for memorized items and productions for generalized items
Convergence Score Average of all values for item-level production similarity for the same items between different learners trained on the same language
Human Label Similarity Item-level production similarity to (other) human learners, averaged across different human learners
True Label Similarity Item-level production similarity to input language
@@ Results
Figure 2: Final Generalization Score achieved by humans (A), GPT-3.5 (B), and recurrent neural networks (C) for each of the input languages. The x-axis shows the structure score of the input languages. Each point corresponds to the generalization score calculated for the entire input language. This score reflects the degree to which learners systematically generalized new labels relative to the labels they learned. For example, generalization score would be high if learners successfully recombines previously used parts, e.g., combining ’muif’ for the shape and ’i’ for the direction into ’muif-i’. Error regions of the regression lines are 95% confidence intervals estimated via bootstrapping. More structure in the input language leads to more systematic generalization for all three learning systems.
@@
Table 2: Generalization examples from neural network and human learners, showing labels generated for unseen scenes. The column GPT-3.5 corresponds to completions generated by the GPT-3.5 model text-davinci-003 via in-context learning, where the training data is provided in context. The examples cover the differently structured input languages from low to high.
Struct. Shape Angle Human RNN GPT-3.5
@@
Strikingly, the results reveal that a higher degree of compositional structure in the input language leads to generalizations that are more systematic (see Figure 2B), closely reflecting the pattern of adult human learners (Figure 2A). Table 2 shows examples of the final productions of humans and large language models during generalization (more examples are provided in Tab. 6 and 7 in the SI).
In addition, we evaluate the production similarity as character-level length-normalized edit distance between the generated labels and labels produced by human participants during generalization. The results show that, given more structured linguistic input, GPT-3.5 also yields productions that are more similar to the productions of human participants, calculated as the average similarity between GPT-3.5’s production and all human productions for the same scene in the same language (Figure 3B). Analogously, Figure 3A shows the similarity of humans to other human learners during generalization.
Figure 3: Final Similarity to Humans during Generalization Final production similarity with (other) human participants during generalization achieved by humans (A), GPT-3.5 (B) and recurrent neural networks (C) for each of the input languages. The x-axis shows the structure score of the input languages. Each point corresponds to the production similarity score (calculated as length-normalized edit distance) between humans’ productions and models’ productions for every item in the language. For example, a recurrent neural network that produced ’muif-a’ for shape 3 moving in direction 360 degrees would have a high production similarity to the majority of human participants who produced ’muif-i’. Error regions of the regression lines show 95% confidence intervals estimated via bootstrapping. More structure in the input language leads to more similarity to human participants for both RNNs and GPT-3.5.
@@
Figure 4: Memorization Error Analysis for human participants (A), GPT-3.5 (B), and recurrent neural networks (C). The error rates are 33.30% for humans, 7.39% for GPT-3.5 via in-context learning, and 13.87% for RNNs after 100 epochs of training. The x-axis shows the structure score of the input language. Each point corresponds to the production similarity score (calculated as length-normalized edit distance) between an erroneously memorized label for a given item and the correct corresponding label as it appears in the input language. For example, ’wangsus’ has a higher similarity with ’wangsuus’ than ’gempt’. Error bands of the regression lines show 95% confidence intervals estimated via bootstrapping. More structure leads to erroneously memorized examples being more similar to the ground truth of the input language.
@@
Figure 5: Learning trajectory of Recurrent Neural Networks’ Memorization and Generalization Performance. More structured languages lead to better and faster reproduction of the input language (A), to better generalization on unknown scenes (C), better agreement with human participants during memorization (B) and generalization (D), and higher convergence between networks (E). (A): Production similarity between labels generated by neural agents and labels of the input language. (B): Production similarity between labels generated by neural agents and labels generated by human participants. (C): Generalization score of labels generated by neural agents for new scenes that were not part of the training data. (D): Production similarity between labels generated by neural agents and labels generated by human participants for unseen scenes. (E): Convergence score measures the similarity between labels generated for unseen scenes by different neural agents. Stars mark the round at which neural agents first exceed the final performance of human participants. Input languages are grouped into 5 bins. Each line is the average of 200 neural agents with a different random initialization. A star marks the epoch at which the RNN agents exceed human performance. Results are cut off for visualization at epoch 60, full results in SI.
@@
The results are shown in Figure 5. Extended results can be found in Figs. 6–11 in the SI. In the following, we present the results for the learning trajectory organized along the two types of tests: memorization and generalization, before presenting the final results of RNNs.
@@
Similarity to Input Languages during Memorization
With sufficient training rounds, all languages can be learned by all neural network agents, reaching a production similarity of at least 0.8 (out of 1) by round 60 (Figure 5A). Structured languages are learned significantly better (LME 1; $\beta=0.045$, $\mathrm{SD}=0.001$, $z=62.865$, $p<0.001$), i. e., they show a higher similarity with the input language. However, this advantage tends to diminish over training rounds (LME 1; $\beta=-0.005$, $\mathrm{SD}<0.001$, $z=-54.978$, $p<0.001$).
Similarity to Humans during Memorization
@@
Similarity to Humans during Generalization
We measure the similarity between the productions of neural network agents and humans for new scenes (Figure 5D), i. e., during generalization. Examples are shown in Table 2. More structure in the input language led to a significantly higher similarity between humans and neural agents (LME 5; $\beta=0.132$, $\mathrm{SD}=0.002$, $z=70.280$, $p<0.001$), which became stronger over rounds ($\beta=0.046$, $\mathrm{SD}<0.001$, $z=344.287$, $p<0.001$).
Convergence between Neural Agents during Generalization
More structured languages lead to better agreement between networks (LME 4; $\beta=0.043$, $\mathrm{SD}=0.001$, $z=49.027$, $p<0.001$), such that, for more structured languages, different neural agents learning the same input language produced more similar labels for new scenes (Figure 5E). This effect became stronger over rounds ($\beta=0.009$, $\mathrm{SD}<0.001$, $z=121.740$, $p<0.001$).
@@
To compare our custom recurrent neural network agents with large language models and with humans, we visualize the relationship between compositional structure of the input language and final generalization performance in Figure 2C. All three learning systems (Humans, RNNs, and GPT-3.5) show the same trend: more compositionality in the input language leads to more systematic generalization.
@@
Moreover, we calculate the average similarity to generalizations of human participants on the same language and item. Comparing the productions during generalization, the results show that a higher degree of structure in the input language leads to more similarity with humans (see Figure 3C). This pattern of compositional structure leading to more human-like generalizations is present in both RNNs’ and GPT-3.5’s generated labels – as well as when comparing humans to other humans (see Figure 3).
@@
Lastly, we visualize the results of the memorization error analysis for recurrent neural networks alongside humans and GPT-3.5 in Figure 4. The pattern is the same for all three different learning systems, be it artificial or biological: more compositional structure leads to errors that are more similar to the true label.
Discussion
@@
Our results show that deep neural networks benefit from more structured linguistic input as humans do and that neural networks’ performance becomes increasingly more human-like when trained on more structured languages. This structure bias can be found in networks’ learning trajectories and even more so in the networks’ generalization behavior, mimicking previous findings with humans. Although all languages can eventually be (almost) perfectly learned, we show that more structured languages are learned better and more similarly to human productions. Deep neural networks and humans produce nearly identical labels when trained on high-structured languages but not when trained on low-structured languages. Moreover, networks that learn more structured languages are significantly better at systematic generalization to new, unseen items, and crucially, their generalizations are significantly more consistent and more human-like. This means that highly systematic grammars allow for better generalization and facilitate greater alignment between different neural agents and between neural agents and humans. We have replicated these results with small recurrent neural networks and with transformer-based large language models, showing that, together with humans, all three learning systems show the same bias in systematic generalization and memorization errors. Thus, our findings strengthen the idea that language models are useful for studying human cognitive mechanisms, complementing the increasing evidence of similarity in language learning between humans and machines [42, 43, 44, 45, 46, 47, 48].
Specifically, we find very similar effects of structure on generalization and on the similarity to humans across all three learning systems. While we find a different slope for humans and RNNs in the memorization error analysis (likely due to RNNs being less impacted by memorization difficulty given sufficient training), the general trend is consistent: for both humans and artificial agents, exposure to more structured languages leads to production errors that are nevertheless more similar to the correct labels (i.e., their errors are less “wrong”).
We assume that the reason for the increased similarity between machines and humans is that the ways to generalize are more transparent in high-structured languages, while there are none or less transparent generalization patterns available in low- and medium-structured languages. This leads both humans and neural networks to a higher production variation in lower structured languages, as different options on how to generalize are equally likely. This point is well supported by results from humans, who indeed show increased convergence between participants when learning higher structured languages [28] Our results thereby demonstrate that what is more transparent for humans is also more transparent for deep neural networks.
Analyzing the learning trajectory of recurrent neural networks, we find that languages with mid and mid-low structures often show an advantage in both memorization and generalization during the early stages of learning. This may be due to the fact that these mid-structured languages trade off full expressiveness with more simplicity (see Tab. 3 in the SI). For example, one of the mid-structured languages includes a marker for “moving on the diagonal”, but does not distinguish the direction of the movement (e.g., center to north-east vs. center to south-west). As a result, the same label is used for two distinct meanings, which is easier to learn in the first place (less variation), but not sufficient to fully differentiate between items and thus harming systematic generalization.
@@
As for implications, our findings first and foremost support the idea that languages’ underlying grammatical structure can be learned directly from (grounded) linguistic input alone [74, 75, 76, 41, 35]. To ensure that the advantage of more structured linguistic input does not stem from the fact that the learning system was already proficient in a different language – i. e., as are pre-trained language models and adult humans – we also also considered models trained from random initialization. Therefore, our results predict that children would also benefit from more systematic compositional structure in the same way adults do – a prediction we are currently testing (preregistration: [77]).
@@
Our findings have further implications for machine learning, where systematic generalization beyond the training distribution (out-of-domain) is of high interest [17, 19, 78, 20, 21]. Systematic in-domain generalization, as studied here, is a critical prerequisite for systematic out-of-domain generalization. Specifically, we show that seeding a learning system with well-structured inputs can improve their ability to systematically generalize to combinations that were not observed during training. Even though our study is based on artificial languages, our findings directly pertain to the natural language processing of real-world languages. To confirm this prediction, we re-analyzed data from Wu et al. [14], who used the Wug Test [79] to test language models’ ability to predict different forms of unfamiliar words in a wide range of natural languages. Indeed, we find that the Wug Test accuracy negatively correlates with the degree of irregularity of the language (Spearman’s $\rho=-0.96$, $p<10^{-15}$; Kendall’s $\tau=-0.86,p<10^{-14}$). This strong negative correlation suggests that natural languages with fewer irregularities, i. e., more consistently structured natural languages, are indeed easier to learn for machines.
@@
Crucially, there is a positive correlation between the degree of linguistic structure and population size [12, 80, 81, 68], with low-resource languages (i. e., languages spoken by smaller communities for which there is only very little training data available) typically having less structured languages. Since our study predicts that such languages are harder to learn for deep neural networks, this results in a double whammy for developing natural language processing systems for small communities’ languages – exacerbating challenges of low-resource language modeling [82]. Interestingly, the benefit of structured input could also explain the importance of highly-structured programming languages in the data mix for training large language models [83].
@@
An interesting direction for future research is to investigate potential differences in the amount of training that a neural network needs compared to humans. Through anchoring our experiments in human data, we were able to directly identify the point during training at which recurrent neural networks equalize with human participants. However, the location of this point depends on various factors such as the amount of data, the number of parameters that are optimized, and the number of optimization steps – which makes it challenging to predict this point in advance. While we have identified this point through analyzing the learning trajectory, our analysis does not depend on it, as all measures including the similarity between humans and machines are calculated based on productions taken at the end of training.
Furthermore, we have chosen to work with an input representation that we deemed easiest to process for each type of learning system. Since the particular way in which agents represent the visual world was not the object of the current study, our rationale here was to provide each learning system with a representation that is easiest or most natural to process. Human participants would have likely had a harder time finding patterns in attribute-value vectors, consisting of 6 numbers, than in short video clips with moving objects. In contrast, operating on raw pixels is expected to introduce more difficulty for machine learning models in terms of disentangling representations [85]. Future work could examine whether neural nets segment visual stimuli in a similar way as humans in grounded language learning.
In conclusion, our findings shed light on the relationship between language and language-learning systems, showing that linguistic structure is crucial for language learnability and systematic generalization in neural networks. Our results suggest that more structured languages are easier to learn, regardless of the learning system: a human, a recurrent neural network, or a large language model. Thus, generalization capabilities are heavily influenced by compositional structure, with both biological and artificial learning systems benefitting from more structured input by facilitating more systematic and transparent generalizations. In future work, we will analyze how this learnability bias for more structure affects neural networks engaged in collaborative communication games, and test how this kind of systematic structure arises in the first place in emergent communication simulations. Moreover, our findings give a clear prediction that children would benefit from more structure in the linguistic input, which we will test by conducting a learnability study with children.
Methods